Be sure to answer all parts.

If 0.050 mol hydrochloric acid (HCl) is mixed with 0.050 mol of sodium hydroxide (NaOH) in a "coffee cup" calorimeter, the temperature of 1.33 × 102 g of the resulting solution increases from 25.3°C to 31.0°C.

Part 1 out of 2 If the specific heat of the solution is 1.00cal/g·°C, calculate the quantity of energy involved in the reaction.

Q = Enter your answer in scientific notation.
Answer #1

Answer – We are given, moles of HCl = 0.050 moles , moles of NaOH = 0.050 moles , mass of solution = 1.33 *102 g , ti = 25.3 oC , tf = 31.0o C

C = 1.00 cal/goC

We know formula for the heat

q = m * C * ∆T

    = 1.33 *102 g * 1.00 cal/goC * (31.0 – 25.3)oC

    = 758.1 cal

we know,

1 cal = 4.184 J

so, 758.1 cal = ? J

= 3171.9 J

= 3.17*103 J

Since heat is evolved it is with negative sign

The quantity of energy involved in the reaction 3.17*103 J